OpenStudy (anonymous):

What is the product of the radical expression? ( sqrt 2+ sqrt 10) (sqrt 2 - sqrt 10)

OpenStudy (anonymous):

\[(\sqrt{2}+ \sqrt{10}) (\sqrt{2}-\sqrt{10})\]

OpenStudy (anonymous):

A. \[-8\sqrt{\sqrt{10}}\] B. \[12\] C. -8 D. \[8\sqrt{2}\]

OpenStudy (anonymous):

\[(\sqrt{2}+\sqrt{10})(\sqrt{2}-\sqrt{10})=\sqrt{4}+\sqrt{20}-\sqrt{20}-\sqrt{100}\]

OpenStudy (anonymous):

that should be easy to simplify

OpenStudy (anonymous):

ugh i cant understand

OpenStudy (anonymous):

can you show me the answer and explain why it is that answer so i can do the next problems easier?

OpenStudy (anonymous):

first you are multiplying each number, doesn't really matter if you do the square root right now after multiplying you can clearly see that 20 - 20 = 0 so that's gone

OpenStudy (anonymous):

now you are left with square root of 4 which is 2 cause 2*2 = 4 and square root of 100 which is 10 cause 10*10 =100 so 2-10 = ....

OpenStudy (anonymous):

8

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

yep